# The Impact of Large Language Models (LLMs) on Code Review Process

### v1

@@ 6.1 Construct Validity
Identifying GPT Use via Explicit Mentions: the use of a single keyword (”GPT”) to identify PRs that potentially involve the assistance from ChatGPT. This keyword-based method may result in false negatives, where GPT assistance was used but not explicitly mentioned, as well as false positives, where the term ”GPT” appears in a different context (e.g., in a discussion). Consequently, this introduces potential bias in the classification of PRs as GPT-assisted and may affect the accuracy of our findings regarding the actual impact of GPT use on the review process. While a semi-automated verification step helps mitigate this issue, future work could benefit from more robust detection methods, such as NLP-based intent analysis.
Definition of GPT as Assistance: A potential threat lies in how the GPT assistance is operationally defined. This study relies on heuristics to label PRs as GPT-assisted or non-assisted, but these heuristics may inadvertently misclassify instances, particularly when contributors do not explicitly document their use of GPT in PR titles, comments, or files. To mitigate this, manual inspections were performed to refine the heuristics and achieve high labeling precision (95%). However, some misclassification may persist.
Measurement of PR Phases: The study measures the duration of PR phases (e.g., review, waiting) using timestamps and events extracted from GitHub. While these metrics provide an objective basis for analysis, they may not fully capture qualitative aspects, such as the complexity of interactions or the nature of GPT’s contributions. In addition, inconsistencies in how timestamps and events are logged across projects could introduce variability. Efforts to mitigate these issues include cross-checking phase definitions and excluding PRs with missing or illogical events and timestamps.
Impact Attribution: Attributing observed reductions in PR phase times to GPT assistance alone is challenging. Other factors, such as task simplicity or team dynamics, could contribute to the observed effects. The use of a linear model helps account for some confounding variables (e.g., project age, number of commits), but residual confounding remains a potential threat. By addressing these threats, the study aims to provide a robust assessment of GPT’s influence on PR phases. However, future research could refine the constructs further, for example, by incorporating direct user feedback on their interactions with GPT or using more sophisticated techniques to distinguish GPT contributions from other factors.
Interactions with GPT: The nature of developer interactions with GPT may not be fully captured by simple time measurements. For example, the impact of GPT on developer behavior (e.g., whether it encourages more efficient coding practices or simply accelerates repetitive tasks) is not directly measured in this study. Future research could incorporate qualitative data, such as developer surveys or interviews, to better understand the cognitive and collaborative effects of GPT assistance on the code review process.
@@ 6.2 Internal Validity
Confounding Variables: The impact of GPT on PR resolution times may be influenced by factors other than GPT assistance, such as the complexity of the code changes, the experience of the developer, or the team’s workflow dynamics. While the study controls for some confounding variables through regression models, residual effects may still persist. Further steps, such as matching or propensity score analysis, could help control for additional confounders that were not included in the model.
Task Categorization: Categorizing the tasks for which GPT is used could be subjective and may vary depending on the developer’s interpretation of GPT’s assistance. If task categorization is inconsistent or unclear, it could skew the analysis and affect the conclusions drawn about the types of assistance most commonly used.
The Role of the Linear Model in Time Reduction: While the study finds a reduction in time for various phases of the code review process when GPT is used, it is important to ensure that this reduction in time is genuinely due to GPT’s assistance and not influenced by other factors (e.g., more efficient developers or less complex PRs). If time reduction is attributed to factors other than GPT, it could undermine the internal validity of the study.
@@ 6.3 External Validity
Platform-Specific Bias: This study analyzes GitHub data, which may have unique characteristics that limit the generalizability of findings to other platforms such as GitLab or Bitbucket. For instance, GitHub as a large open-source community and due to its widespread usage may introduce dynamics that are not presented in smaller or private repositories. Future research could include multiple platforms to better assess the broader applicability of these findings.
Repository Selection: The data included in this study were selected based on a threshold of 10 stars, indicating a relatively active user base. While this ensures that the results reflect commonly used, publicly available code, it also introduces selection bias. The results may not generalize well to less popular or private repositories, where developer practices may differ. Expanding the dataset to include a wider range of repositories, including those with lower visibility or fewer contributors, would help assess the external validity of the findings.
Task Variability: PRs vary widely in terms of content, including code size, task complexity, and the nature of the proposed changes. Although we attempted to control for these factors, the diversity in PRs introduces a level of variability that could influence the generalizability of the results. To address this, future studies could consider stratified sampling or clustering based on PR characteristics to assess the impact of GPT more accurately across different types of tasks.
@@ 7 Conclusion
This study highlights the potential of generative AI tools, particularly large language models (LLMs) such as ChatGPT, to enhance collaborative software development. Through an analysis of their application in GitHub pull requests (PRs), we observe substantial improvements across key phases of the PR workflow, most notably within the review phase, which exhibits a 33% reduction in review duration and an 87% decrease in waiting time for acceptance. These results underscore the capacity of GPT to improve code review efficiency and support diverse tasks, including code enhancement, bug resolution, and documentation generation, with the most significant gains occurring during the review process. While the contributions of GPT are clearly beneficial in certain phases, its limited influence during the submission and post-acceptance phases points to areas for further development and optimization. Future research should investigate how GPT can be more effectively integrated across the entire PR lifecycle and evaluate its impact on long-term productivity. As LLMs continue to advance, they are poised to become essential components of software engineering, promoting greater efficiency, collaboration, and code quality. By harnessing the capabilities of these tools, both practitioners and researchers can shape the evolution of collaborative development workflows and drive innovation in the field of software engineering.
Appendix A
In the following, we provide a detailed list of the patterns found in project name, PR title, PR body, and modified file to identify PRs using GPT as assistance, along with their definitions and intended matching purpose.
P1: Pattern for Detecting GPT-Related Identifiers. This regex identifies phrases containing ”chatgpt,” ”gpt,” or ”openai,” ensuring that they are preceded by descriptive words. It prevents ”chat” or ”gpt” from appearing immediately after the separator, enforcing a structured naming convention for AI-related terms. Example: project chatgpt_123.
[…]
P2: Pattern for Detecting GPT-Related Keywords in URL. Matches URLs containing ”chatgpt”,”gpt”, or ”openai” while ensuring a valid structure. It supports both HTTP and HTTPS protocols, optional ”www” subdomains, and common top-level domains (TLDs). The pattern also captures query parameters and paths where AI-related keywords appear, making it ideal for filtering relevant web links. Example:https://my-site.ai/chatgpt-version
[…]
P3:Pattern in Detecting with GPT-related Keywords in Path and URL. matches file pathsthat include specific AI-related keywords like ”chatgpt,” ”gpt,” or ”openai,” while also supporting common file path structures. It handles both forward and backslashes, valid path components, and file extensions. Example: /home/user/gpt_data.csv
[…]
|(chatgpt|gpt|openai))+\.[a-z]{2,3})
P4: Pattern in Detecting GPT Keywords in File Paths. matches strings containing the keywords ”chatgpt,” ”gpt,” or ”openai” within file paths or URLs. It handles optional spaces, separators (like hyphens and underscores), and supports valid URL components (letters, digits, and special characters). It’s designed to identify AI-related content in file paths, query parameters, or directories while allowing flexible formatting. Example: chatgpt@example.com, chatbot@openai.com
[…]
P5: Pattern for Matching ChatGPT References in Phrases . matches phrases like ”like in the chatgpt” or ”like the one from chatgpt” at the end of a string. It efficiently captures variations in the phrasing, ensuring that the string ends with a reference to ”chatgpt” while supporting both expressions.
[…]
P6: Pattern for Matching ”Spam” Followed by AI Keywords. matches the word ”spam” followed by a space and one of the AI-related keywords (”chatgpt,” ”gpt,” or ”openai”) at the end of a string. It is optimized for simplicity and flexibility, ensuring accurate detection of such phrases.
[…]
P7: Pattern for Matching AI Keywords and Code-Related Terms. matches strings containing AI-related keywords (like ”chatgpt,” ”gpt,” ”openai,” or ”azureai”) followed by code-related terms such as ”label,” ”function,” ”class,” ”model,” or ”service.” It handles both singular and plural forms efficiently, making it ideal for detecting AI and code-related concepts in text or code.
[…]
(label|class|functions?|constructors?|
calls?|formatter|demos?|sources?|
connectors?|plugins?|endpoints?|
models?|apis?|gateways?|services?))
P8: Pattern for Matching AI Keywords with Numeric Suffixes. matches occurrences of AI-related keywords (”gpt”, ”chatgpt”, ”openai”) followed by optional spaces, hyphens, or underscores and a numeric suffix (up to 6 digits or decimals). It is designed to efficiently identify AI keywords with numeric versions or identifiers, and works for both single and repeated instances in the string.
[…]
P9: Pattern for Matching GRUB GPT References. matches instances of ”grub_gpt” or ”GRUB_GPT” in strings, with support for optional spaces, hyphens, underscores, digits, and periods preceding the keywords. It identifies these specific references in various contexts, including file paths, configuration strings, or any similar patterns.
[…]
P10: Pattern for Matching AI Keywords at the End of a String. matches strings that end with one of the AI-related keywords (”chatgpt”, ”gpt”, or ”openai”), allowing for optional spaces, URL-safe characters, and hyphens before the keyword. It is ideal for identifying these specific keywords in paths, URLs, or other contexts where they appear at the end of a string.
[…]
(chatgpt|gpt|openai)\$
P11: Pattern for Matching Disk Types and Partition Formats. matches common disk types and partition formats, including ”disk guid,” ”mbr,” ”qemu,” ”kvm,” and ”gpt partition(s).” It is optimized for detecting variations in case and optional spaces or plurals.
(disk guid|mbr|MBR|qemu|QEMU|kvm|
[…]
P12: Pattern for Matching ChatGPT Requests with Action Verbs. matches requests directed to ”ChatGPT” or ”GPT” using various action verbs like ”say,” ”write,” ”suggest,” ”think,” etc., allowing for optional spaces or hyphens between ”chat” and ”gpt.” It’s designed to identify common user queries to AI models, including instructions and requests for explanations or suggestions.
[…]
(say|get|write|suggest|
tell|think|explain|
[…]
P13: Pattern for Matching OpenAI Chat Links. matches URLs from https://chat.openai.com, including optional /share/ paths and any query parameters. It efficiently captures both the base URL and potential query strings, ensuring flexible matching of OpenAI chat links.
https:\/\/chat\.openai\.com(\/share\/?)
[…]
P14: Pattern for Matching Examples Referencing GPT and Related Models. matches phrases like ”some example by GPT,” ”some example by ChatGPT,” or similar phrases involving AI models like OpenAI, Chat-GPT, or Coderabbit. It captures variations in spacing, hyphenation, and model names while ensuring flexibility in the format of such references.
[…]
(gpt|chatgpt|openai|chat\-gpt|
[…]
P15: Pattern for Matching GPT Versions Followed by ”Flawless”. matches strings like ”gpt-123 flawless,” where ”gpt-” is followed by numeric values and any combination of alphanumeric characters and spaces, concluding with the word ”flawless.”
[…]
P16: Pattern for Matching GPT Requests to ”Make Some Suggestion”. matches strings like ”gpt-123 make some suggestion,” where ”gpt-” is followed by digits and any alphanumeric characters or spaces, with the phrase ”make some suggestion” at the end. It accounts for optional spaces in the phrase for flexibility.
[…]
P17: Pattern for Matching GPT Requests to ”Walk Me”. matches the phrase ”gpt walk me,” allowing for optional spaces between ”gpt,” ”walk,” and ”me,” making it flexible for user input variations.
[…]
P18: Pattern for Matching Actions Performed with GPT and Related Models. captures actions like ”ask,” ”create,” ”generate,” and more, followed by prepositions like ”by,” ”with,” or ”from,” and then identifies associations with GPT and related models (e.g., ”gpt,” ”chatgpt,” ”openai,” etc.). It’s ideal for recognizing specific action-based queries related to AI.
[…]
generate|document|investigate|advise|
suggest|modify|solve|support|update|
summary|response|change|power|polish|
utilize|comment|write|improve|collaborate|
[…]
@@
Antonio Collante: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing Original Draft, Writing - Review and Editing, Visualization Samuel Abedu: Conceptualization, Software, Data Curation, Writing - Review and Editing SayedHassan Khatoonabadi: Conceptualization, Methodology, Validation, Writing - Review and Editing, Project Administration Ahmad Abdellatif: Conceptualization, Validation, Writing - Review and Editing Ebube Alor: Data Curation Emad Shihab: Conceptualization, Supervision, Project Administration, Funding Acquisition

### v2

@@ 6.1 Construct Validity
Construct validity concerns whether the study measurements accurately capture the concept of GPT assistance and its effects on the PR review process. In this study, GPT use was identified through explicit mentions of the keyword “GPT” in PRs, which may introduce false negatives (when GPT assistance is used but not explicitly mentioned) and false positives (when “GPT” appears in unrelated contexts). This limitation could bias the classification of PRs as GPT-assisted and affect the precision of the findings. Although a semi-automated verification step mitigated this issue, future work could employ more robust NLP-based intent analysis for detection. Another threat arises from how GPT assistance is operationally defined. The study uses heuristic rules to label PRs as GPT-assisted or non-assisted, which may lead to misclassifications when contributors do not document GPT use explicitly. To address this, manual inspections were conducted to refine the heuristics, achieving high labeling precision (95%), and substantial inter-rater agreement was confirmed using Cohen’s Kappa, with discrepancies resolved by consensus among authors. The measurement of PR phases (e.g., review and waiting) using GitHub timestamps and events provides an objective foundation but may overlook qualitative aspects such as interaction complexity or the nature of the contributions by GPT. Inconsistencies in event logging across projects were mitigated by cross-checking phase definitions and excluding anomalous data. Furthermore, attributing observed reductions in PR phase durations solely to GPT assistance remains challenging, as other factors, such as task simplicity or team dynamics, could also influence outcomes. Although linear modeling accounted for some confounders (e.g., project age, number of commits), residual confounding may persist. Finally, developer interactions with GPT are not fully captured by temporal metrics; aspects such as behavioral or cognitive effects require further qualitative exploration. Future studies could integrate developer surveys or interviews to deepen understanding of how GPT shapes collaboration and review practices.
@@ 6.2 Internal Validity
Internal validity addresses the extent to which the observed effects can be causally attributed to GPT assistance rather than other influencing factors. Several potential confounding variables, such as code change complexity, developer experience, and team workflow dynamics, can have affected PR resolution time independently of GPT use. Although the study controls for some of these confounders using regression models, residual effects could still persist. Future research could strengthen control through methods such as matching or propensity score analysis to account for unobserved variables. Another concern relates to task categorization since developers may interpret or record their use of GPT differently, leading to inconsistencies in how tasks are classified. Such subjectivity could skew results and obscure patterns regarding the types of assistance most frequently used. Finally, while the study observes reduced time across various code review phases for GPT-assisted PRs, it remains crucial to verify that these reductions are genuinely attributable to the contributions of GPT rather than from external factors such as more efficient developers or inherently simpler tasks. Ensuring that time savings reflect true GPT influence is essential to maintaining the internal validity of the findings.
@@ 6.3 External Validity
External validity concerns the extent to which the findings of this study can be generalized beyond the specific context analyzed. Since the dataset is derived exclusively from GitHub, the results may be influenced by platform-specific characteristics, such as community norms or review workflows, that differ from those on other platforms like GitLab or Bitbucket. The study also focuses on repositories with at least ten stars, ensuring a baseline level of activity but potentially introducing selection bias, as less popular or private repositories may exhibit different development and review practices. Additionally, the wide variability among PRs (in terms of code size, complexity, and task type) poses challenges to generalization, as these factors can affect both the use and impact of GPT assistance. To enhance external validity, future research could incorporate data from multiple platforms, include a broader range of repositories, and apply stratified sampling or clustering techniques to better capture task diversity and contextual differences
@@ 7 Conclusion
This study highlights the potential of generative AI tools, particularly LLMs such as ChatGPT, to enhance collaborative software development. Through an analysis of their application in GitHub PRs, we observe substantial improvements across key phases of the PR workflow, most notably within the review phase, which exhibits a 66.7% reduction in review time and an 87.5% decrease in waiting time for acceptance. These results show the capacity of GPT to improve code review efficiency and support diverse tasks, including code enhancement, bug resolution, and documentation generation, with the most significant gains occurring during the review process. While the contributions of GPT are clearly beneficial in certain phases, its limited influence during the submission and post-acceptance phases points to areas for further development and optimization. Future research should investigate how GPT can be more effectively integrated across the entire PR lifecycle and evaluate its impact on long-term productivity. As LLMs continue to advance, they are poised to become essential components of software engineering, promoting greater efficiency, collaboration, and code quality. By harnessing the capabilities of these tools, both practitioners and researchers can shape the evolution of collaborative development workflows and drive innovation in the field of software engineering.
@@
Antonio Collante: Conceptualization, Methodology, Software, Investigation, Data Curation, Writing Original Draft, Writing - Review and Editing, Visualization Samuel Abedu: Conceptualization, Software, Data Curation, Writing - Review and Editing SayedHassan Khatoonabadi: Conceptualization, Methodology, Validation, Writing - Review and Editing, Project Administration Ahmad Abdellatif: Conceptualization, Validation, Writing - Review and Editing Ebube Alor: Data Curation Emad Shihab: Conceptualization, Supervision, Project Administration, Funding Acquisition
@@ Acknowledgements
Appendix A Tables
In the following, we provide the GPT PRs during a Round 1, the accuracy across rounds after applying heuristics, and a summary of the GPT Assisted PR Dataset by Status.
Table 5: GPT PRs during a Round of Sampling Analysis
Filter Assistance Non Assistance
Project name 4 21
PR title 6 19
Modified files 23 58
PR body 1 24
Total 34 122
Table 6: Accuracy Improvement across Rounds after Applying Heuristics
Heuristic Round 1 (%) Round 2 (%) Round 3 (%)
H1 82 98 98
H2 98 97 100
H3 96 97 97
H4 66 95 97
Note: Each round corresponds to one iteration of refinement, where heuristics were introduced to correct common misclassifications.
Table 7: Summary: GPT Assisted PR Dataset by Status
Status Assistance
Open 54
Closed 179
Merged 1367
Total 1600